OpenStudy (anonymous):

Calculus: Find the area of the surface obtained by rotating the curve y=2x (2 to 5) about the x-axis.

OpenStudy (anonymous):

\[I set \it up like this: \int\limits_{0}^{1}2\pi(2x)\sqrt(5)\]

OpenStudy (anonymous):

The setup is almost correct, the limits should be \(\displaystyle\int_2^5\cdots\)